OpenStudy (anonymous):

Which of the graphs below represents the equation 3x - 2y = -12?

OpenStudy (anonymous):

wheres the graphs?

OpenStudy (anonymous):

to find this get the equation in y=mx+b form m being the slope and b being y intercept 3x-2y=-12 subtract both sides by 3 -2y=-3x-12 now divide by -2 \[y=\frac{ 3 }{ 2 }x+6\]

OpenStudy (anonymous):

to graph start on the y axis go up 6 next for the slope go up 3 over to the right 2 should look similar to that

OpenStudy (anonymous):

which one has a positive slope (line leans to the right) of 3/2